# Efficient Individually Rational Recommender System under Stochastic Order

### v1

@@ Abstract
With the rise of online applications, recommender systems (RSs) often encounter constraints in balancing exploration and exploitation. Such constraints arise when exploration is carried out by agents whose individual utility should be balanced with overall welfare. Recent work [5] suggests that recommendations should be individually rational. Specifically, if agents have a default arm they would use, relying on the RS should yield each agent at least the reward of the default arm, conditioned on the knowledge available to the RS. Under this individual rationality constraint, striking a balance between exploration and exploitation becomes a complex planning problem. We assume a stochastic order of the rewards (e.g., Bernoulli, unit-variance Gaussian, etc.), and derive an approximately optimal algorithm. Our technique is based on an auxiliary Goal Markov Decision Process problem that is of independent interest. Additionally, we present an incentive-compatible version of our algorithm.
@@ 1 Introduction
Multi-armed bandits (MABs) represent a fundamental problem in online learning. In this setting, the goal is to maximize total rewards by balancing the exploitation of known high-reward arms with the exploration of potentially better ones. With the rise of online applications, MAB scenarios often involve human elements, where arms or rounds represent individuals whose needs must be considered. For instance, several lines of work consider content providers as the arms [17, 9, 19], requiring algorithms to incentivize the creation of high-quality content. Therefore, besides balancing exploration and exploitation, the algorithm must also address fairness [20, 24], long-term welfare [9], strategic behavior [12, 15], etc.
@@
A recent line of work, pioneered by the seminal paper of Kremer et al. [23], considers MAB as a mechanism design problem [29, 30]. Recommender systems suggest actions to agents without forcing their choices. Since agents are selfish and may avoid high-risk, high-reward arms, there is a need to incentivize exploration to ensure accurate reward estimates for all arms. Previous works [25, 27, 14] achieve incentive compatibility by leveraging information asymmetry: While each agent possesses some prior information, the mechanism accumulates additional information by observing the actions and rewards of previous agents. Despite addressing agents’ incentives, this approach has a major drawback: Since exploring agents could have gotten better rewards had they not used the mechanism, they might see the recommendations of the IC mechanism as betraying their trust.
To address this issue, Bahar et al. [5] introduce the concept of individual rationality (IR) in MAB mechanisms. IR, a celebrated concept in mechanism design [33], ensures that each recommendation’s expected reward, based on the mechanism’s knowledge, meets or exceeds the reward from the agent’s default arm. Relying on the mechanism’s knowledge contrasts with information asymmetry and limits the mechanism’s ability to learn through exploration. Consequently, IR and IC mechanisms must adopt a careful exploration policy, maximizing their ability to explore risky arms and thereby maximizing welfare. Such an exploration policy should be randomized, offering portfolios that mix low-risk, low-reward and high-risk, high-reward arms. Due to their complex structure, prior work [5] has presented solutions with inefficient runtime, particularly exponential in the number of arms. This paper is, to the best of our knowledge, the first to propose a runtime-efficient IR and IC MAB mechanism for non-trivial reward distributions.
@@ 1.1 Our Contribution
We consider a MAB setting with a Bayesian prior and static rewards; that is, the reward of each arm is initially unknown and is realized only once. [^1] Although this assumption is limiting, it allows us to focus on the novel IR viewpoint proposed in this paper. A similar approach is taken by several early works on incentivizing exploration [4, 14, 5]. Despite the limitation, the problem remains technically non-trivial. Some arms may be risky, having negative expected rewards but also offering high potential rewards. Additionally, we assume that there is a default arm that agents would pick if they did not use the algorithm. In each round, the algorithm picks a distribution over the arms, termed a portfolio. The algorithm then samples and selects an arm based on the portfolio’s weights. We aim to maximize social welfare, defined as the sum of rewards.
Without any further constraints, maximizing welfare is straightforward: Try each arm once, then repeatedly choose the best one. Due to our static rewards assumption, one round of exploration per arm is enough to reveal the highest reward. However, we restrict the algorithm to select only individually rational (IR) portfolios. A portfolio is IR if its expected reward is at least as good as that of the default arm conditioned on the information available to the algorithm (no information asymmetry). Since rewards are realized only once, careful planning is crucial. Arms with a high expected value extend the set of IR portfolios, enabling the exploration of a priori inferior arms.
Our main technical contribution is an asymptotically optimal IR algorithm under a stochastic order assumption on rewards (see Definition 2.2). Stochastic order is common in practical scenarios, e.g., Bernoulli rewards or Gaussian reward distributions with a common variance. We introduce an auxiliary Goal Markov Decision Process (GMDP) [8]. In our GMDP, the action sets are convex polytopes of all IR portfolios that mix a priori superior arms (better than the default arm) with a priori inferior arms. Despite the complex action space, we reveal a simple index-based optimal policy, akin to the seminal result of Weitzman [35] for Pandora’s box [16, 2, 11, 10]. We then use the optimal GMDP policy to develop IREGB (see Algorithm 2), an IR and asymptotically optimal algorithm. Informally stated,
Theorem 1.1 (Informal version of Corollary 4.1).
Given a problem instance with a stochastic order over rewards, IREGB has a runtime of $O(K\log K)$, where $K$ is the number of arms, and asymptotically achieves optimal welfare among all IR algorithms.
We analyze the algorithm’s convergence rate and demonstrate that it is fast in expectation. Additionally, we assume agents are strategic and aim to design an IC and IR mechanism. To achieve this, we apply the hidden exploration technique of Mansour et al. [25] to modify IREGB, making it both IR and IC.
@@ 1.2 Related Work
Individual rationality (IR) can be viewed as a safety constraint, connecting to various strands of research on safe exploration. For instance, Bandits with knapsack [3] address a MAB problem with a global budget, aiming to maximize total rewards before exhausting resources. Another approach focuses on stage-wise safety, ensuring that regret performance remains above a threshold set by a baseline strategy in each round [21, 36]. Notably, in these lines of work, constraints apply to cumulative resource consumption or reward throughout the algorithm’s run. Conversely, Amani et al. [1] apply a reward constraint in each round. Their work leaves the set of safe decisions uncertain due to the inherent uncertainty in the learning process, aiming to minimize regret while learning the safe decision set. Similar to Amani et al. [1], we also consider a stage-wise constraint. It is worth mentioning that safety has been studied beyond MABs [34, 28]. For example, Moldovan and Abbeel [28] introduce an algorithm enabling safe exploration in Markov Decision Processes to avoid fatal absorbing states.
Closely related to our work is the research on incentivizing exploration, pioneered by Kremer et al. [22], motivated by using a bandit-like setting for recommendations. Since users are selfish, algorithms must incentivize exploration (see [32] for an introduction and overview). Subsequent works consider regret minimization [25, 27], heterogeneous agents [13, 18], social networks [4, 7], and extensions to Markov Decision Processes [31]. Similarly to several previous works [5, 6, 14, 22], we assume each arm is associated with a fixed value, which is initially unknown and sampled from a known distribution. Given this, our problem becomes the careful planning of safe exploration.
Bahar et al. [5] is the work most closely related to ours within the incentivizing exploration line of research. We inherit their model with minor differences and also aim to develop an optimal individually rational mechanism. Bahar et al. [5] assume rewards are integers in a bounded set of size $H$, construct an auxiliary GMDP model, and devise a dynamic programming-based planning policy with a runtime of $O(2^{K}K^{2}H^{2})$. We adopt the same Goal MDP approach (Section 3) but aim to reduce the exponential dependence on $K$. We achieve a significant time improvement of $O(K\log K)$ for rewards with stochastic order, a special yet common case. Importantly, the techniques we use, such as the Equivalence Lemma (Lemma 3.1), are novel and were not examined by Bahar et al. [5].
@@ 2 Problem Statement
Definition 2.1 (Bayesian Incentive Compatibility).
@@
$\operatorname{\mathbb{E}}(X(a_{l})-X(a_{i})\mid m^{t}(\mathcal{I}_{t})=a_{l})\geq 0.$ (1)
@@
Individual Rationality
In this paper, we focus on ex-ante individually rational recommendations, abbreviated as IR. A portfolio $\boldsymbol{p}^{t}$ is IR at time $t$ if, given the information the mechanism acquired up to round $t$, its expected reward is greater or equal to zero, which is the reward of the default arm. Formally,
Definition 2.2 (Ex-ante Individual Rationality).
A portfolio $\boldsymbol{p}$ is individually rational (IR) with respect to $\mathcal{I}$ if
@@
We stress that Definition 2.2 implies only ex-ante IR, as Nature may realize an arm with a negative expectation; nevertheless, we shall continue referring to it as IR. If a mechanism selects IR portfolios for every possible information $\mathcal{I}$, we say the mechanism is IR. By focusing on IR mechanisms, we ensure that every agent will receive at least the reward of the default arm in expectation, independent of the other agents. The definitions of IR and IC are orthogonal and significantly differ in the events they condition on. The IC Inequality (1) conditions on the agent’s information, which includes the prior information and the recommendation of the mechanism. In contrast, the IR Inequality (2) conditions on all the information the mechanism has acquired. Forbidding the mechanism from exploring arms with a negative expected value trivializes the problem as it prevents exploring any such arms. The IR constraint is a compromise: It allows learning about a priori inferior arms, but only when they are mixed with a priori superior arms.
@@
When agents follows the mechanism, we denote the social welfare achieved by a mechanism ALG by $SW_{T}({\small\textnormal{{ALG}}})=\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}r^{t}\right]$. The global objective of the mechanism is to maximize social welfare, subject to selecting IR portfolios in every round. If agents are strategic, we narrow our attention to BIC mechanisms only, which justifies using the same formulation of social welfare.
@@ 2.1 Example
Example 2.1.
Let $K=4$, and let $X(a_{1})\sim N(2,1)$, $X(a_{2})\sim N(1,1)$, $X(a_{3})\sim N(-1,1)$, and $X(a_{4})\sim N(-2,1)$, where $N(\mu,\sigma^{2})$ denotes the normal distribution with a mean $\mu$ and a variance $\sigma^{2}$. Notice that $\textnormal{pos}(A)=\{a_{1},a_{2}\}$ and $\textnormal{neg}(A)=\{a_{3},a_{4}\}$. In the first round, the mechanism can select, for example, a portfolio that comprises $a_{1}$ w.p. 1. It can also select the portfolio that mixes $a_{1}$ w.p. $\frac{1}{2}$ with $a_{2}$ w.p. $\frac{1}{2}$, and infinitely many other IR portfolios.
To demonstrate the technical difficulty of maximizing social welfare, assume that $X(a_{1}),X(a_{2})<0$ while $X(a_{3}),X(a_{4})>0$ (this information is not available to the mechanism). Indeed, this occurs with positive non-negligible probability under the distributional assumptions of this example. Consider the case where the mechanism selects the portfolio $a_{1}$ w.p. 1 in the first round and $a_{2}$ w.p. 1 in the second round. Notice that these portfolios are IR: Given the information $\mathcal{I}_{1}$ in the first round, selecting $a_{1}$ w.p. 1 is IR, and regardless of $x(a_{1})$, selecting $a_{2}$ w.p. 1 in the second round is also IR. After observing that $a_{1}$ and $a_{2}$ have negative rewards (since we momentarily assume $X(a_{1})<0$ and $X(a_{2})<0$), the only IR portfolio for the subsequent rounds is to select the default arm $a_{0}$ w.p. 1.
However, there are much better ways to act in the first round. Arms with positive expected rewards are essential for enabling the exploration of arms with negative expected rewards. Consider $\boldsymbol{p}^{1}$ such that $\boldsymbol{p}^{1}(a_{1})=\frac{1}{3},\boldsymbol{p}^{1}(a_{3})=\frac{2}{3}$ and $\boldsymbol{p}^{1}(a_{2})=\boldsymbol{p}^{1}(a_{4})=0$. Notice that $\boldsymbol{p}^{1}$ is IR w.r.t. $\mathcal{I}_{1}$, since $\boldsymbol{p}^{1}(a_{1})\mu(a_{1})+\boldsymbol{p}^{1}(a_{3})\mu(a_{3})=0$. If the mechanism selects $\boldsymbol{p}^{1}$ in the first round, it will already discover the positive reward of arm $a_{3}$ in the first round w.p. $\frac{2}{3}$. It is thus evident that selecting $\boldsymbol{p}^{1}$ in the first round results in higher social welfare than selecting $a_{1}$ w.p. 1.

### v2

@@ Abstract
With the rise of online applications, recommender systems (RSs) often encounter constraints in balancing exploration and exploitation. Such constraints arise when exploration is carried out by agents whose utility must be taken into account when optimizing overall welfare. Recent work suggests that recommendations should be mechanism-informed individually rational (MIR) [6]. Specifically, if agents have a default arm they would use, relying on the RS should yield each agent at least the reward of the default arm, conditioned on the information available to the RS. Under the MIR constraint, striking a balance between exploration and exploitation becomes a complex planning problem. To that end, Bahar et al. [6] propose an approximately optimal yet inefficient planning algorithm that runs in $O(2^{K}K^{2}H^{2})$, where $K$ is the number of arms and $H$ is the size of the support of the reward distributions. In this paper, we make a significant improvement for a special yet practical case, removing both the dependence on $H$ and the exponential dependence on $K$. We assume a stochastic order of the rewards (e.g., Gaussian with unit variance, Bernoulli, etc.), and devise an asymptotically optimal algorithm with a runtime of $O(K\log K)$. Our technique is based on formulating a Goal Markov Decision Process (GMDP), establishing an optimal dynamic programming procedure, and then unveiling its crux—fleshing out a simple index-based structure that facilitates efficient computation. Additionally, we present an incentive-compatible version of our algorithm.
Keywords:
E
xplore-exploit, Recommender Systems, Individual Rationality, Incentive compatibility, Mechanism Design.
@@ 1 Introduction
Multi-armed bandits (MABs) represent a fundamental problem in online learning. In this framework, the goal is to maximize total rewards by balancing the exploitation of known high-reward arms with the exploration of potentially better ones. With the rise of online applications, MAB scenarios often involve human-facing settings, where arms or rounds represent agents whose needs must be considered. For instance, several lines of work consider content providers as the arms [17, 8, 19], requiring algorithms to incentivize the creation of high-quality content. In other works, beyond the exploration-exploitation dilemma, algorithms must also address fairness [20, 24], long-term welfare [9, 28], and strategic behavior [12, 15], among other concerns.
@@
A recent line of work initiated by Kremer et al. [23] considers MAB as a mechanism design problem [30, 31]. In such settings, a recommender system suggests actions to agents without enforcing its choices. As agents are self-interested and may avoid high-risk, high-reward arms, there is a need to incentivize exploration to ensure accurate reward estimates for all arms. Previous works [25, 27, 14] achieve incentive compatibility (IC) by leveraging information asymmetry: While each agent possesses some prior information, mechanisms accumulate additional information by observing the actions and rewards of previous agents. Although IC typically serves as a benchmark for a mechanism’s trustworthiness, as it guarantees that agents maximize their expected utility by following the mechanism’s recommendations, mechanisms that employ information asymmetry still suffer from an inherent trust issue. Indeed, agents who receive recommendations that are deliberately suboptimal (i.e., exploration) may feel misled.
One intuitive remedy is to modify the mechanism beyond mere IC by promising that an agent is never worse off by following a recommendation. This brings us to the classical notion of Individual Rationality (IR), or participation constraints [34]. IR requires that the expected utility from following the recommendation be at least as high as that of the agent’s default action—that is, the action they would select if the mechanism did not exist. Yet the standard IR definition inherits the same information-asymmetry pitfall: Because the mechanism holds more information than any single agent, it can propose actions that look rational from the agent’s perspective but are sub-optimal relative to the mechanism’s more comprehensive information.
To address this issue, Bahar et al. [6] propose a stronger guarantee, Mechanism-informed Individual Rationality (MIR). [^1] Bahar et al. [6] label this guarantee simply “IR.” To avoid confusion with the conventional IR constraint, we henceforth call it mechanism-informed IR. MIR requires that every recommendation’s expected reward, based on all the information available to the mechanism, meet or exceed the agent’s default-arm reward ex-ante. This stronger guarantee is crucial because it closes the information gap, treating the agent as if they possessed the mechanism’s accumulated information. Despite the attractiveness of MIR, relying on the mechanism’s accumulated information limits the ability to learn through exploration. Consequently, MIR mechanisms must adopt a careful exploration policy, maximizing their ability to explore risky arms while still satisfying the MIR guarantee by mixing low-risk and high-risk arms.
Due to this challenge, prior work [6] faces significant practical limitations. Indeed, the runtime of the MIR mechanism of Bahar et al. [6] is $O(2^{K}K^{2}H^{2})$, where $K$ is the number of arms and $H$ is the support size of the (discrete) reward distributions. This high runtime complexity severely limits the mechanism’s practicality, particularly when dealing with many arms or continuous reward distributions. For the latter, discretization approaches are either computationally infeasible or yield suboptimal rewards.
In this paper, we revisit MIR constraints in Bayesian MABs with static rewards. For priors that are stochastically ordered, we design IREGB: An index-based algorithm that is MIR by construction and is asymptotically optimal among all MIR algorithms. Crucially, IREGB runs in $O(K\log K)$ time—reducing the exponential $2^{K}K^{2}$ dependence on the number of arms down to near-linear and removing the $H^{2}$ term entirely, so it works with continuous reward distributions. Finally, we use IREGB as a black-box to develop a MIR and IC algorithm.
@@ 1.1 Our Contribution
We consider a MAB setting with a Bayesian prior and static rewards; that is, the reward of each arm is initially unknown and is realized only once. [^2] This assumption allows us to focus on the novel mechanism-informed IR viewpoint proposed in this paper. A similar approach is taken by several early works on incentivizing exploration [4, 6], and some have strictly stronger assumptions, i.e., $\{-1,1\}$ rewards [14]. Despite the limitation, our problem remains technically non-trivial. Some arms may be risky, having negative expected rewards but also offering high potential rewards. Additionally, we assume that there is a default arm that agents would pick if they did not use the algorithm. In each round, the algorithm picks a distribution over the arms, termed a portfolio. The algorithm then samples and selects an arm based on the portfolio’s weights. We aim to maximize social welfare, defined as the sum of rewards.
Without any further constraints, maximizing welfare is straightforward: Pull each arm once, then repeatedly choose the best one. Due to our static rewards assumption, one round of exploration per arm is enough to reveal the highest reward. However, we restrict the algorithm to select only MIR portfolios. A portfolio is MIR if its expected reward is at least as good as that of the default arm conditioned on the information available to the algorithm (no information asymmetry). Since rewards are realized only once, careful planning is crucial. Arms with a high expected value extend the set of MIR portfolios, enabling the exploration of a priori inferior arms.
Our main technical contribution is an asymptotically optimal MIR algorithm under a stochastic order assumption on rewards (see Definition 2). Stochastic order is common in practical scenarios, e.g., Bernoulli rewards or Gaussian reward distributions with a common variance. We introduce an auxiliary Goal Markov Decision Process (GMDP) [7]. In our GMDP, the action sets are convex polytopes of all MIR portfolios that mix a priori superior arms (better than the default arm) with a priori inferior arms. Despite the complex action space, we reveal a simple index-based optimal policy, akin to the seminal result of Weitzman [36] for Pandora’s box [16, 2, 11, 10]. We then use the optimal GMDP policy to develop IREGB (see Algorithm 1), a MIR and asymptotically optimal algorithm. Informally stated,
Theorem 1.1 (Informal version of Corollary 1)
Given a problem instance with a stochastic order over rewards, IREGB has a runtime of $O(K\log K)$, where $K$ is the number of arms, and asymptotically achieves optimal welfare among all MIR algorithms.
We analyze the algorithm’s convergence rate and demonstrate that it is fast in expectation. Additionally, we assume agents are strategic and aim to design an IC and MIR mechanism. To achieve this, we apply the hidden exploration technique of Mansour et al. [25] to modify IREGB, making it both MIR and IC.
@@ 1.2 Related Work
Mechanism-informed individual rationality can be viewed as a safety constraint, connecting to various strands of research on safe exploration. For instance, Bandits with knapsack [3] address a MAB problem with a global budget, aiming to maximize total rewards before exhausting resources. Another approach focuses on stage-wise safety, ensuring that regret performance remains above a threshold set by a baseline strategy in each round [21, 37]. Notably, in these lines of work, constraints apply to cumulative resource consumption or reward throughout the algorithm’s run. Conversely, Amani et al. [1] apply a reward constraint in each round. Their work leaves the set of safe decisions uncertain due to the inherent uncertainty in the learning process, aiming to minimize regret while learning the safe decision set. Similar to Amani et al. [1], we also consider a stage-wise constraint. It is worth mentioning that safety has been studied beyond MABs [35, 29]. For example, Moldovan and Abbeel [29] introduce an algorithm enabling safe exploration in Markov Decision Processes to avoid fatal absorbing states.
Closely related to our work is the research on incentivizing exploration, pioneered by Kremer et al. [22], motivated by using a bandit-like setting for recommendations. Since users are selfish, algorithms must incentivize exploration (see [33] for an introduction and overview). Subsequent works consider regret minimization [25, 27], heterogeneous agents [13, 18], social networks [4, 5], and extensions to Markov Decision Processes [32]. Similarly to several previous works [6, 5, 14, 22], we assume each arm is associated with a fixed value, which is initially unknown and sampled from a known distribution. Given this, our problem becomes the careful planning of MIR exploration.
Bahar et al. [6] is the work most closely related to ours within the incentivizing exploration line of research. We inherit their model and also aim to develop an optimal MIR mechanism. Bahar et al. [6] assume rewards are integers in a bounded set of size $H$, construct an auxiliary GMDP model, and devise a dynamic programming-based planning policy with a runtime of $O(2^{K}K^{2}H^{2})$. We adopt the same GMDP approach (Section 3) but aim to reduce the exponential dependence on $K$ and remove the dependence on $H$. We achieve a significant improvement of $O(K\log K)$ for rewards with stochastic order, a special yet common case. Importantly, the techniques we use, such as the Equivalence Lemma (Lemma 1), are novel and were not examined by Bahar et al. [6].
@@ 2 Problem Statement
Definition 1 (Bayesian Incentive Compatibility)
@@
$\operatorname{\mathbb{E}}[X(a_{l})-X(a_{i})\mid m^{t}(\mathcal{I}_{t})=a_{l}]\geq 0.$ (1)
@@
Mechanism-Informed Individual Rationality
In this paper, we focus on ex-ante mechanism-informed individually rational recommendations, abbreviated as MIR. A portfolio $\boldsymbol{p}^{t}$ is MIR at time $t$ if, given the information the mechanism acquired up to round $t$, its expected reward is greater than or equal to zero, which is the reward of the default arm. Formally,
Definition 2 (Ex-ante Mechanism-Informed Individual Rationality (MIR))
A portfolio $\boldsymbol{p}$ is MIR with respect to information $\mathcal{I}$ if
@@
We stress that Definition 2 implies only ex-ante MIR, as Nature may realize an arm with a negative expectation; nevertheless, we shall continue referring to it as MIR. If a mechanism selects MIR portfolios for every possible information $\mathcal{I}$, we say the mechanism is MIR. By focusing on MIR mechanisms, we ensure that every agent will receive at least the reward of the default arm in expectation, independent of the other agents. The definitions of MIR and BIC are orthogonal and significantly differ in the events they condition on. The BIC Inequality (1) conditions on the agent’s information, which includes the prior information and the recommendation of the mechanism. In contrast, the MIR Inequality (2) conditions on all the information the mechanism has acquired. Thereby, MIR could be viewed as a standard for how much and in what way the system may leverage its informational advantage.
We note that forbidding the mechanism from exploring arms with a negative expected reward (equivalent to demanding ex-post MIR) trivializes the problem as it prevents exploring any such arms. The MIR constraint is a compromise: It allows learning about a priori inferior arms, but only when they are mixed with a priori superior arms.
@@
When agents follow the mechanism, we denote the social welfare achieved by a mechanism ALG by $SW_{T}({\small\textnormal{{ALG}}})=\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}r^{t}\right]$. The global objective of the mechanism is to maximize social welfare, subject to selecting MIR portfolios in every round. If agents are strategic, we narrow our attention to BIC mechanisms only, which justifies using the same formulation of social welfare.
@@ 2.1 Example
Example 1
Let $K=4$, and let $X(a_{1})\sim N(2,1)$, $X(a_{2})\sim N(1,1)$, $X(a_{3})\sim N(-1,1)$, and $X(a_{4})\sim N(-2,1)$, where $N(\mu,\sigma^{2})$ denotes the normal distribution with a mean $\mu$ and a variance $\sigma^{2}$. Notice that $\textnormal{pos}(A)=\{a_{1},a_{2}\}$ and $\textnormal{neg}(A)=\{a_{3},a_{4}\}$. In the first round, the mechanism can select, for example, a portfolio that comprises $a_{1}$ w.p. 1. It can also select the portfolio that mixes $a_{1}$ w.p. $\frac{1}{2}$ with $a_{2}$ w.p. $\frac{1}{2}$, and infinitely many other MIR portfolios.
To demonstrate the technical difficulty of maximizing social welfare, assume that $X(a_{1}),X(a_{2})<0$ and $X(a_{3}),X(a_{4})>0$ (this information is not available to the mechanism). Indeed, this occurs with positive non-negligible probability under the distributional assumptions of this example. Consider the case where the mechanism selects the portfolio $a_{1}$ w.p. 1 in the first round and $a_{2}$ w.p. 1 in the second round. Notice that these portfolios are MIR: Given the information $\mathcal{I}_{1}$ in the first round, selecting $a_{1}$ w.p. 1 is MIR, and regardless of $x(a_{1})$, selecting $a_{2}$ w.p. 1 in the second round is also MIR. After observing that $a_{1}$ and $a_{2}$ have negative rewards (since we momentarily assume $X(a_{1})<0$ and $X(a_{2})<0$), the only MIR portfolio in subsequent rounds is selecting the default arm $a_{0}$ w.p. 1.
However, there are better ways to act in the first round. Arms with positive expected rewards are essential for enabling the exploration of arms with negative expected rewards. Consider $\boldsymbol{p}^{1}$ such that $\boldsymbol{p}^{1}(a_{1})=\frac{1}{3},\boldsymbol{p}^{1}(a_{3})=\frac{2}{3}$ and $\boldsymbol{p}^{1}(a_{2})=\boldsymbol{p}^{1}(a_{4})=0$. Notice that $\boldsymbol{p}^{1}$ is MIR w.r.t. $\mathcal{I}_{1}$, since $\boldsymbol{p}^{1}(a_{1})\mu(a_{1})+\boldsymbol{p}^{1}(a_{3})\mu(a_{3})=0$. If the mechanism selects $\boldsymbol{p}^{1}$ in the first round, it will already discover the positive reward of arm $a_{3}$ in the first round w.p. $\frac{2}{3}$. It is thus evident that selecting $\boldsymbol{p}^{1}$ in the first round results in higher social welfare than selecting $a_{1}$ w.p. 1.